# The MASSIVE Survey. XX. A Triaxial Stellar Dynamical Measurement of the Supermassive Black Hole Mass and Intrinsic Galaxy Shape of Giant Radio Galaxy NGC 315

### v1

@@ 2.2.1 Central stellar kinematics
We observed the central $5^{\prime\prime}{\times}7^{\prime\prime}$ of NGC 315 using the two-slit IFS mode of Gemini Multi-Object Spectrograph (GMOS) on the Gemini North Telescope with 1000 hexagonal lenslets (each with a projected diameter of $0\farcs 2$). Ten science exposures of 1200 seconds each were obtained, totaling 3.3 hours of on-source and simultaneous observations of a $5^{\prime\prime}{\times}3\farcs 5$ region of sky offset by ${\sim}1^{\prime}$ from the galaxy. The R400400400400-G5305530553055305 grating with the CaT filter was used to cover the wavelength range $7800{-}9330$. The median seeing was $0\farcs 6$ FWHM, and the median spectral resolution (determined from arc lamp lines for each lenslet) was $2.5$ FWHM. Details of the data processing procedure are described in MASSIVE Paper XIII (Ene et al., 2019).
@@ 3.1 The TriOS Code and Galaxy Models
We use the TriOS code (Quenneville et al., 2021, 2022) to compute triaxial orbit models of NGC 315. This code integrates a large number of stellar orbits that span the allowed phase space and computes the LOSVDs for a wide range of galaxy model parameters. The galaxy is assumed to have three mass components: a central black hole of mass $M_{\mathrm{BH}}$, a stellar component with a mass-to-light ratio $M^{*}/L$, and a dark matter halo with a density profile $\rho(r)=\rho_{0}/[(r/r_{s})^{\gamma}\left(1+r/r_{s}\right)^{3-\gamma}]$, where $r_{s}$ is a scale radius (Navarro et al., 1996). We set $\gamma=0$ so that the profile has a finite central density $\rho_{0}$ and a flattened central density distribution similar to that of the stars. With our data, $r_{s}$ and $\rho_{0}$ are often quite degenerate so we choose to parameterize the halo with a single parameter, $M_{15}$, defined to be the dark matter mass enclosed within $15$ kpc with a fixed scale radius $r_{s}=15$ kpc. A similar strategy was used in orbit modeling of other MASSIVE galaxies (e.g., Liepold et al. 2020; Quenneville et al. 2022; Pilawa et al. 2022).
@@
[…]
@@
Mass and Shape Parameters of NGC 315
Galaxy Property (units) Value
@@ 3.2 Best-Fit Triaxial Model
We use the model selection scheme described in our recent work (e.g., Pilawa et al. 2022; Liepold et al. 2023, 2025) to determine the mass and shape parameters that best match the kinematic and photometric data of NGC 315. This scheme, in brief, involves generating candidate galaxy models with Latin hypercube sampling (McKay et al., 1979), approximating the resulting likelihood landscapes with Gaussian process regression, and sampling posteriors on our parameters via dynamic nested sampling (Speagle, 2020). About ${\sim}1750$ galaxy models are used to obtain the final posteriors on the six parameters shown in Figure 4. An additional ${\sim}3000$ models covering wider ranges of parameters are used in the initial exploration of the 6D likelihood surface, and ${\sim}10,000$ more models are used in various tests to ensure the surface in the low $\chi^{2}$ region is mapped out accurately.
@@
A summary of the best-fit parameters for NGC 315 is listed in Table 1. The kinematic moments predicted by the model are compared with observed values in Figure 3. The total $\chi^{2}$ is $1611.4$ spread over $2314$ kinematic constraints ($8$ moments for 248 GMOS bins and $6$ moments for 55 Mitchell bins). A naive estimate of the reduced $\chi^{2}$ would be $1611.4/(2314-6)=0.698$ assuming 6 degrees of freedom (DOF). But we caution that DOF is nontrivial to estimate for nonlinear problems such as here. In a study using simulated stellar kinematics that mimic data in the MASSIVE survey (Pilawa et al., 2024), we have investigated a “generalized” measure of DOF for 6-parameter models and found it to be ${\sim 200}$ instead of the canonical value of 6, thereby raising the reduced $\chi^{2}$ by ${\sim}40$% in that study. While a similar calculation would have to be performed to estimate this DOF measure for NGC 315, we expect the reduced $\chi^{2}$ to be raised as well.
3.2.1 Black Hole Mass
@@
To assess how NGC 315 and its SMBH fit in the population of local galaxies with dynamically inferred $M_{\mathrm{BH}}$, we examine its location on the well-studied $M_{\mathrm{BH}}{-}\sigma$ and $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relations. NGC 315’s luminosity-weighted velocity dispersion within $R_{e}$ is found to be $\sigma_{e}=341\text{ km s}^{-1}$ based on the same Mitchell IFS data in this paper (Veale et al., 2017b). At this $\sigma_{e}$, the mean $M_{\mathrm{BH}}{-}\sigma$ relations of McConnell & Ma (2013) and Saglia et al. (2016) predict $M_{\mathrm{BH}}=4.2\times 10^{9}M_{\odot}$ and $3.3\times 10^{9}M_{\odot}$, respectively, $40\%$ (0.15 dex) and $10\%$ (0.04 dex) larger than our dynamically measured $M_{\mathrm{BH}}$. However, our $M_{\mathrm{BH}}$ is within the intrinsic scatter of both relations, $0.38$ dex.
@@
To place NGC 315 on the $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relation, we use the total stellar mass from our best-fitting triaxial model, $M_{*}=1.5\times 10^{12}M_{\odot}$. At this bulge mass, the mean relations of McConnell & Ma (2013) and Saglia et al. (2016) predict $M_{\mathrm{BH}}=5.3\times 10^{9}M_{\odot}$ and $4.4\times 10^{9}M_{\odot}$, respectively, $77$% (0.25 dex) and $47$% (0.17 dex) larger than our measured $M_{\mathrm{BH}}$. But again, our dynamical $M_{\mathrm{BH}}$ is within the intrinsic scatter of both $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relations, $0.34$ dex.
@@ 3.2.2 Intrinsic 3D Galaxy Shape
The preferred axis ratios of $p=0.854$ and $q=0.833$ yield a triaxiality parameter of $T=\left(1-p^{2}\right)/\left(1-q^{2}\right)=0.89{\pm}0.02$, where the limits of $T=0$ and 1 correspond to oblate axisymmetry and prolate axisymmetry, respectively. Thus, NGC 315 is triaxial and highly prolate. In comparison, five other massive elliptical galaxies for which we have performed triaxial orbit modeling thus far all have smaller $T$. Four of them are oblate ($T\lesssim 0.5$): $T=0.33{\pm}0.06$ for NGC 1453 (Quenneville et al., 2021), $T=0.39{\pm}0.04$ for NGC 2693 (Pilawa et al., 2022), $T=0.35{\pm}0.03$ for Holmberg 15A (Liepold et al., 2025), and $T=0.31{\pm}0.05$ for NGC 57 (Pilawa et al. 2025, in prep). M87, on the other hand, is slightly prolate with $T=0.65{\pm}0.02$ (Liepold et al., 2023).
@@ 3.2.3 Stellar Orbital Structure
The composition of the major types of orbits in the best-fit model for NGC 315 is plotted as a function of radius in the upper panel of Figure 5. The two types of tube orbits have a fixed sense of rotation, with short-axis tubes having an angular momentum component along the minor axis that does not change sign, and similarly, long-axis tubes having an angular momentum component along the major axis that does not change sign. For the box orbits, all three components of the angular momentum change sign, leaving no sense of rotation for this orbit type. The relative contribution of these three orbit types determines the velocity structure of the galaxy.
Figure 5: Composition of the three major orbit types (upper panel) and velocity anisotropy (lower panel) as a function of radius in the best-fit triaxial galaxy model of NGC 315. The majority of the orbital weights are in long-axis tube orbits, consistent with the prolateness of the galaxy. Box orbits and short-axis tube orbits contribute about $25\%$ and $20\%$ of the remaining orbital weights, respectively. The velocity anisotropy parameters $\beta\equiv 1-\sigma_{t}^{2}/\sigma_{r}^{2}$ (in spherical coordinates) and $\beta_{z}=1-\sigma^{2}_{z}/\sigma^{2}_{R}$ (in cylindrical coordinates) indicate the stellar orbits are mildly tangentially anisotropic at small radii and are increasingly radially anisotropic at larger radii.
Figure 5 shows that beyond the central ${\sim}200\text{ pc}$, the fraction of long-axis tubes, short-axis tubes, and box orbits is nearly a constant of ${\sim}55\%$, ${\sim}20\%$, and ${\sim}25\%$, respectively. The predominance of long-axis tubes over short-axis tubes reflects the prolateness of NGC 315. This is opposite to the orbital compositions in oblate triaxial galaxies with $T<0.5$, e.g., about $60{-}70$% of orbits are short-axis tubes while ${\sim}20$% are long-axis tubes in NGC 1453 and NGC 2693 (Quenneville et al., 2021; Pilawa et al., 2022).

### v2

@@ 2.2.1 Central stellar kinematics
We observed the central $5^{\prime\prime}{\times}7^{\prime\prime}$ of NGC 315 using the two-slit IFS mode of Gemini Multi-Object Spectrograph (GMOS) on the Gemini North Telescope with 1000 hexagonal lenslets (each with a projected diameter of $0\farcs 2$). Ten science exposures of 1200 seconds each were obtained, totaling 3.3 hours of on-source and simultaneous observations of a $5^{\prime\prime}{\times}3\farcs 5$ region of sky offset by ${\sim}1^{\prime}$ from the galaxy. The R400400400400-G5305530553055305 grating with the CaT filter was used to cover the wavelength range $7800{-}9330$. The median spectral resolution (determined from arc lamp lines for each lenslet) was $2.5$ FWHM. Details of the data processing procedure are described in MASSIVE Paper XIII (Ene et al., 2019).
@@ 3.1 The TriOS Code and Galaxy Models
We use the TriOS code (Van Den Bosch et al., 2008; Quenneville et al., 2021, 2022) to compute triaxial orbit models of NGC 315. This code integrates a large number of stellar orbits that span the allowed phase space and computes the LOSVDs for a wide range of galaxy model parameters. The galaxy is assumed to have three mass components: a central black hole of mass $M_{\mathrm{BH}}$, a stellar component with a mass-to-light ratio $M^{*}/L$, and a dark matter halo with a density profile $\rho(r)=\rho_{0}/[(r/r_{s})^{\gamma}\left(1+r/r_{s}\right)^{3-\gamma}]$, where $r_{s}$ is a scale radius (Navarro et al., 1996). We set $\gamma=0$ so that the profile has a finite central density $\rho_{0}$ and a flattened central density distribution similar to that of the stars. With our data, $r_{s}$ and $\rho_{0}$ are often quite degenerate so we choose to parameterize the halo with a single parameter, $M_{15}$, defined to be the dark matter mass enclosed within $15$ kpc with a fixed scale radius $r_{s}=15$ kpc. A similar strategy was used in orbit modeling of other MASSIVE galaxies (e.g., Liepold et al. 2020; Quenneville et al. 2022; Pilawa et al. 2022).
@@
In total, each galaxy model has $6$ free parameters – $M_{\mathrm{BH}}$, $M^{*}/L$, $M_{15}$, and three shape parameters – to be constrained by the kinematic and photometric data. For each model, we use the same procedures for phase space sampling and orbit integrations as in our earlier work (e.g., Section 4.1 of Liepold et al. 2023). We integrate the trajectories of stars to build a library of 437,400 stellar orbits. A dithering factor of $N_{\rm dither}=3$ is used, where the properties of $N_{\rm dither}^{3}=27$ orbits with similar initial conditions are averaged, effectively giving them a single shared orbital weight. We determine the resulting 16,200 independent orbital weights using the kinematic measurements with non-negative least squares (Lawson & Hanson, 1995), under the additional constraint that both the projected mass within each aperture and the 3D mass distribution in coarse bins are consistent within 1% of the MGE. The GMOS and Mitchell PSFs are taken to be single, circularly symmetric Gaussians with FWHM of $0\farcs 38$ and $1\farcs 2$, respectively.
@@
Galaxy Parameters of NGC 315 from Triaxial Orbit Modeling
Galaxy Property (units) Inferred Value
@@ 3.2 Best-Fit Triaxial Model
We use the model selection scheme described in our recent work (e.g., Pilawa et al. 2022; Liepold et al. 2023, 2025) to determine the mass and shape parameters that best match the kinematic and photometric data of NGC 315. This scheme, in brief, involves generating candidate galaxy models with Latin hypercube sampling (McKay et al., 1979), approximating the resulting likelihood landscapes with Gaussian process regression, and sampling posteriors on our parameters via dynamic nested sampling (Speagle, 2020). The relative likelihood for each model is computed from
$-2\ln\mathcal{L}(\boldsymbol{\mu})=\chi^{2}_{\textrm{kin}}(\boldsymbol{\mu})=\sum_{j}^{N_{\textrm{bin}}}\sum_{i}^{N_{\textrm{GH}}}\frac{(h_{ij}(\boldsymbol{\mu})-h_{ij,\rm{data}})^{2}}{\Delta h_{ij,\rm{data}}^{2}}\,,$ (1)
where $h_{ij}$ is the $i$-th Gauss-Hermite moment of the stellar LOSVDs in the $j$-th spatial bin, $\Delta h_{ij}$ is the measurement uncertainty, and $\boldsymbol{\mu}$ is the set of six parameters describing the galaxy’s potential. To obtain 1D confidence intervals for each galaxy parameter, we take the marginalized distribution for each parameter to be the distribution of (weighted) sample points over the desired parameters, assuming uncertainties in the distribution function (e.g., orbital weights) are neglected.
About ${\sim}1750$ galaxy models are used to obtain the final posteriors on the six parameters shown in Figure 4. An additional ${\sim}3000$ models covering wider ranges of parameters are used in the initial exploration of the 6D likelihood surface, and ${\sim}10000$ more models are used in various tests to ensure that the surface in the low $\chi^{2}$ region is mapped out accurately.
@@
A summary of the best-fit parameters for NGC 315 is listed in Table 1. The kinematic moments predicted by the model are compared with observed values in Figure 3. The total $\chi^{2}$ is $1611.4$ spread over $2314$ kinematic constraints ($8$ moments for 248 GMOS bins and $6$ moments for 55 Mitchell bins). A naive estimate of the reduced $\chi^{2}$ would be $1611.4/(2314-6)=0.698$ assuming 6 degrees of freedom (DOF). But we caution that DOF is nontrivial to estimate for nonlinear problems such as here. In a study using simulated stellar kinematics that mimic data in the MASSIVE survey (Pilawa et al., 2024), we have investigated a “generalized” measure of DOF in the form of a penalty term added to the likelihood measure (Ye, 1998; Lipka & Thomas, 2021) and found it to be ${\sim 200}$ instead of the canonical value of 6, thereby raising the reduced $\chi^{2}$ by ${\sim}40$% in that study. While a similar calculation would have to be performed to estimate this alternative DOF measure for NGC 315, we expect the reduced $\chi^{2}$ to be raised as well. Using this generalized measure to estimate the effective DOF in the models also helps address a related concern that the conventional (unpenalized) likelihood method does not marginalize over the high dimensional space of possible orbital weights, leading to reliable $M_{\mathrm{BH}}$ but with “overly pessimistic error estimates” (Magorrian, 2006). In Pilawa et al. (2024), we find that the parameter search and inference routines used in this paper and other work by our group (e.g., Quenneville et al. 2022; Pilawa et al. 2022; Liepold et al. 2023; Pilawa et al. 2024; Liepold et al. 2025) are able to reliably recover the mass and shape parameters associated with synthetic galaxy kinematic data without any additional penalty terms.
A handful model predictions in Figure 3 lie outside the 68% uncertainties of the data points, e.g., the $h_{4}$ and $h_{6}$ GMOS moments within $0\farcs 6$ are under-predicted (on average) by ${\sim}1.05\sigma$ and ${\sim}1.11\sigma$, respectively, and Mitchell’s velocity dispersion and $h_{6}$ between $5^{\prime\prime}$ and $10^{\prime\prime}$ are over- and under-predicted by ${\sim}1.1\sigma$ and ${\sim}1.6\sigma$, respectively. These systematic local deviations occur in $\lesssim 2\%$ of the 2314 kinematic constraints. When excluding these deviant constraints from the $\chi^{2}$ used for our parameter inference, we find an insignificant change in the inferred $M_{\mathrm{BH}}$.
3.2.1 Black Hole Mass, Stellar Mass, and Dark Matter Mass
@@
To assess how NGC 315 and its SMBH fit in the population of local galaxies with dynamically inferred $M_{\mathrm{BH}}$, we examine its location on the well-studied $M_{\mathrm{BH}}{-}\sigma$ and $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relations. NGC 315’s luminosity-weighted velocity dispersion within $R_{e}$ is found to be $\sigma_{e}=341\text{ km s}^{-1}$ based on the same Mitchell IFS data in this paper (Veale et al., 2017b). At this $\sigma_{e}$, the mean full-sample $M_{\mathrm{BH}}{-}\sigma$ relations of McConnell & Ma (2013) and Saglia et al. (2016) predict $M_{\mathrm{BH}}=4.2\times 10^{9}M_{\odot}$ and $3.3\times 10^{9}M_{\odot}$, respectively, $40\%$ (0.15 dex) and $10\%$ (0.04 dex) larger than our dynamically measured $M_{\mathrm{BH}}$. However, our $M_{\mathrm{BH}}$ is within the intrinsic scatter of both relations, $0.38$ dex.
@@
To place NGC 315 on the $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relation, we use the total stellar mass from our best-fitting triaxial model, $M_{*}=1.5\times 10^{12}M_{\odot}$. At this bulge mass, the mean full-sample relations of McConnell & Ma (2013) and Saglia et al. (2016) predict $M_{\mathrm{BH}}=5.3\times 10^{9}M_{\odot}$ and $4.4\times 10^{9}M_{\odot}$, respectively, $77$% (0.25 dex) and $47$% (0.17 dex) larger than our measured $M_{\mathrm{BH}}$. But again, our dynamical $M_{\mathrm{BH}}$ is within the intrinsic scatter of both $M_{\mathrm{BH}}{-}M_{\rm bulge}$ relations, $0.34$ dex.
In our preferred model, the black hole has a gravitational sphere of influence (SOI) of $r_{\text{SOI}}=0\farcs 81=0.26$ kpc with the definition of $M^{*}(<r_{\text{SOI}})=M_{\mathrm{BH}}$, and $r_{\text{SOI}}=1\farcs 1=0.35$ kpc with the definition of $M^{*}(<r_{\text{SOI}})=2M_{\mathrm{BH}}$.
Within the effective radius $R_{e}=25\farcs 7$ (${\sim}8.3$ kpc) of NGC 315 (Quenneville et al., 2024), the dark matter halo constitutes ${\sim}47\%$ of the total mass of the galaxy. This value is broadly consistent with the rising dark matter fraction with increasing galaxy stellar mass reported for lower-mass early-type galaxies (e.g., Fig. 10 of Cappellari et al. 2013; Fig. 16 of Santucci et al. 2022), where the dark matter fraction (at $1\,R_{e}$) reaches ${\sim}40\%$ at the highest masses ($M^{\star}{\sim}5\times 10^{11}M_{\odot}$) in their samples.
@@ 3.2.2 Intrinsic 3D Galaxy Shape
The preferred axis ratios of $p=0.854$ and $q=0.833$ yield a triaxiality parameter of $T=\left(1-p^{2}\right)/\left(1-q^{2}\right)=0.89{\pm}0.02$, where the limits of $T=0$ and 1 correspond to oblate axisymmetry and prolate axisymmetry, respectively. In comparison, five other massive elliptical galaxies for which we have performed triaxial orbit modeling thus far all have smaller $T$. Four of them are oblate ($T\lesssim 0.5$): $T=0.33{\pm}0.06$ for NGC 1453 (Quenneville et al., 2021), $T=0.39{\pm}0.04$ for NGC 2693 (Pilawa et al., 2022), $T=0.35{\pm}0.03$ for Holmberg 15A (Liepold et al., 2025), and $T=0.31{\pm}0.05$ for NGC 57 (Pilawa et al. 2025, in prep). M87, on the other hand, is slightly prolate with $T=0.65{\pm}0.02$ (Liepold et al., 2023).
@@ 3.2.3 Stellar Orbital Structure
The composition of the major types of orbits in the best-fit model for NGC 315 is plotted as a function of radius in the upper panel of Figure 5. The two types of tube orbits have a fixed sense of rotation, with short- and long-axis tubes having angular momentum components along the intrinsic short- and long-axis, respectively, which do not change sign. For the box orbits, all three components of the angular momentum change sign, leaving no sense of rotation for this orbit type. The relative contribution of these three orbit types determines the velocity structure of the galaxy.
Figure 5: Composition of the three major orbit types (upper panel) and velocity anisotropy (lower panel) as a function of radius in the best-fit triaxial galaxy model of NGC 315. The majority of the orbital weights are in long-axis tube orbits, consistent with the prolateness of the galaxy. Short-axis tube orbits and box orbits both contribute about $20\%-25$% to the remaining orbital weights.
The velocity anisotropy parameters $\beta\equiv 1-\sigma_{t}^{2}/\sigma_{r}^{2}$ (in spherical coordinates) and $\beta_{z}=1-\sigma^{2}_{z}/\sigma^{2}_{R}$ (in cylindrical coordinates) indicate the stellar orbits are mildly tangentially anisotropic at small radii and are increasingly radially anisotropic at larger radii.
Figure 5 shows that $45\%-60\%$ of the mass in the orbits are from long-axis tubes, while short-axis tubes and box orbits contribute comparable amounts to the remaining orbital weight. The predominance of long-axis tubes over short-axis tubes reflects the prolateness of NGC 315. This is opposite to the orbital compositions in oblate triaxial galaxies with $T<0.5$, e.g., about $60{-}70$% of orbits are short-axis tubes while ${\sim}20$% are long-axis tubes in NGC 1453 and NGC 2693 (Quenneville et al., 2021; Pilawa et al., 2022).
Further insight into the relative short- and long-axis tube orbits can be gained from the relationship between the parameter $T$ and the locations in the start space from which the tube orbits are generated (using the scheme of Schwarzschild 1993). Quenneville et al. (2021) shows that the location of the focal curve demarcating the regions in start space populated by long- and short- axis tubes (their Fig. 1) is specified by an angle $\eta$, where $\tan\eta\approx\sqrt{T/(1-T)}$. For $T=0.89$, we have $\eta\approx 71^{\circ}$, indicating the long- and short-axis tube fraction in the start space is $79$% and $21$%, respectively, corresponding to a ratio of about 4 to 1. The actual orbital fractions for the best-fit galaxy model shown in Figure 5 are computed from these base orbits with the corresponding orbital weights, yielding a comparable long-to-short tube ratio of roughly $3$ to $1$. [^4] The preferred viewing angles (in Table 1) are close to the intermediate axis ($\theta$ = $\phi$ = $90^{\circ}$). In this limit, the freedom for the deprojection extends roughly along the intermediate axis, corresponding to different values of $T$ (see Equations (7) and (8) of Quenneville et al. 2022). Within the viewing-angle parameterization of the deprojection, the third viewing angle $\psi$ then sets $T$.